OpenStudy (anonymous):

if the average of 3,4 and x is -2 what is the value of x

OpenStudy (anonymous):

start by thinking of how you find an average of a group of number. You add them all up and then divide by the number of numbers there are, right? so lets base this off of that. \[\frac{ 3+4+x }{ 3}\]=-2 so lets simplify what we have so far \[\frac{ 7+x }{ 3 }=-2 multiply by 3 to get the bottom gone and solve the rest like normal equation 7+x==6 :)